# Cohomology of Lie algebroids over Algebraic spaces

### v7

@@ 5.2. Hochschild hypercohomology of universal enveloping algebroid
In general case, where $\mathcal{L}$ is locally free $\mathcal{O}_{X}$-module of infinite rank (quasicoherent sheaf), there exist a filtered ordered set $J$ as well as an inductive system of free $\mathcal{O}_{X}(U_{x})$-modules $\{{\mathcal{L}(U_{x})}_{j}\mid j\in J\}$ such that
@@
for each special open set $U_{x}$ (using results from [19]). Since both $\mathbb{H}H^{\bullet}$ (which is isomorphic to the derived functor $Cotor^{\bullet}$) and the functor $\mathcal{S}$ commute with inductive limits over a filtered ordered set, we get the result using the result follows from the locally free $\mathcal{O}_{X}$-module of finite rank case (local counterpart is described in [19]). ∎
@@
The HKR morphism is not a ring isomorphism, but composing the HKR-morphism together with the Todd genus provides a ring isomorphism (moreover it provides a canonical Gerstenhaber algebra isomorphism). It is required in the study of formality (or quantization) for Lie algebroids [7].
@@
In the special case of $\mathcal{L}=\mathcal{T}_{X}$ when $X$ is non-singular (or smooth), we get isomorphism of graded vector spaces
@@
((((Noncommutative analogue of HKR Theorem)))) The free Lie algebroid $\mathcal{P}_{X}$ can be identified with sheaf of regular noncommutative vector fields on $X$ and we have seen already that the sections of its universal enveloping algebra $\mathbb{D}_{X}$ are noncommutative differential operators on $X$ ((((see Remark 2.12)))). Thus, in this case from the HKR theorem we get a canonical isomorphism
@@ 5.3. Hochschild hypercohomology of jet algebroid
A canonical flat $\mathcal{L}$-connection, called Grothendieck connection is given by
@@
where $\nabla$ is the Grothendieck connections (Note 5.11), sheafification of the presheaf of canonical left $\mathcal{L}(U)$-connection $U\mapsto\nabla_{U}$ on $\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))$. It is basically the sheafification of the presheaves of cochain complex
@@
where $\bar{K}^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)^{*}=\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))\otimes_{\mathcal{O}_{X}(U)}\wedge^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)^{*}$. The unit of $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ as an $\mathcal{O}_{X}$-algebra provides the morphism from $\mathcal{O}_{X}$ to $\bar{\mathscr{K}}^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*}$.
@@
Since the unit $1_{U}\in\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))$ is given by the counit $\epsilon_{U}$ of $\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U))$, the induced differential is exactly the Lie-Rinehart coboundary $d_{U}$ (4), on an open set $U\subset X$. These isomorphisms are compatible with restrictions.
Now consider the canonical presheaves from the above complexes associated with each open sets, which are isomorphic on each special open sets and compatible with the natural restrictions. Then considering there sheafifications and using the above isomorphisms, we get isomorphism between the complexes of sheaves, described in (14).
@@
for each $U_{x}$ associated with $x\in X$, where $P_{U_{x}}$ is given by the cannonical projections $pr_{1}:\mathcal{J}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))\cong\widehat{S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})^{*}}\rightarrow\mathcal{L}(U_{x})^{*}$ and the anti-symmetrization map $Alt_{U_{x}}:\wedge^{\bullet}_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})^{*}\rightarrow\mathcal{J}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))^{\otimes^{\bullet}}$ for every special open set $U_{x}$ of $X$ (for smooth manifold we get these quasi isomorphisms for each open sets). Sheafification of the associated presheaves provides the quasi-isomorphism between the cobar complex $Cob^{\bullet}(\mathscr{J}(\mathcal{O}_{X},\mathcal{L}))$ and the dual Koszul-Rinehart complex $\bar{\mathscr{K}}^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*}$.
@@
In [5, 6], a dual version of the (twisted) HKR theorem for Lie algebroids [7] (with rich algebraic structures) is considered in-order to study a notion of precalculus upto homotopy in the Lie algebroids context.
@@
All the previous discussions are applicable to Lie algebroids over noetherian separated schemes (or schemes of finite types) over a field of characteristic zero.

### v8

@@ 5.2. Hochschild hypercohomology of universal enveloping algebroid
In general case, where $\mathcal{L}$ is locally free $\mathcal{O}_{X}$-module of infinite rank (quasicoherent sheaf), there exists a filtered ordered set $J$ as well as an inductive system of free $\mathcal{O}_{X}(U_{x})$-modules $\{{\mathcal{L}(U_{x})}_{j}\mid j\in J\}$ such that
@@
for each special open set $U_{x}$ (3.5) (using results from [19]). Since both $\mathbb{H}H^{\bullet}$ (which is isomorphic to the derived functor $Cotor^{\bullet}$) and the functor $\mathcal{S}$ commute with inductive limits over a filtered ordered set, we get the result using the result follows from the locally free $\mathcal{O}_{X}$-module of finite rank case (local counterpart is described in [19]). ∎
@@
The HKR morphism is not a ring isomorphism, but composing it together with the Todd genus provides a ring isomorphism (moreover it provides a canonical Gerstenhaber algebra isomorphism). It is required in the study of formality (or quantization) for Lie algebroids over a ringed site [7].
@@
In the special case of $\mathcal{L}=\mathcal{T}_{X}$ when $X$ is non-singular or smooth ((((see Note 2.2)))), we get isomorphism of graded vector spaces
@@
((((Noncommutative analogue of the HKR Theorem)))) The free Lie algebroid $\mathcal{P}_{X}$ can be identified with sheaf of regular noncommutative vector fields on $X$ and we have seen already that the sections of its universal enveloping algebra $\mathbb{D}_{X}$ are noncommutative differential operators on $X$ ((((see Remark 2.12)))). Thus, in this case from the HKR theorem we get a canonical isomorphism
@@ 5.3. Hochschild hypercohomology of jet algebroid
A canonical flat $\mathcal{L}$-connection on $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$, called Grothendieck connection is given by
@@
where $\nabla$ is the Grothendieck connection (Note 5.12), sheafification of the presheaf of canonical left $\mathcal{L}(U)$-connection $U\mapsto\nabla^{U}$ on $\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))$. It is basically the sheafification of the presheaves of cochain complex
@@
where $\bar{K}^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)^{*}=(\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))\otimes_{\mathcal{O}_{X}(U)}\wedge^{\bullet}_{\mathcal{O}_{X}(U)}\mathcal{L}(U)^{*},\nabla^{U})$. The unit of $\mathscr{J}(\mathcal{O}_{X},\mathcal{L})$ as an $\mathcal{O}_{X}$-algebra provides the morphism from $\mathcal{O}_{X}$ to $\bar{\mathscr{K}}^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*}$.
@@
Since the unit $1_{U}\in\mathcal{J}(\mathcal{O}_{X}(U),\mathcal{L}(U))$ is given by the counit $\epsilon_{U}$ of $\mathcal{U}(\mathcal{O}_{X}(U),\mathcal{L}(U))$, the induced differential is exactly the Lie-Rinehart coboundary $d_{U}$ (4), on an open set $U\subset X$. These isomorphisms are compatible with restrictions. Now consider the canonical presheaves from the above complexes associated with each open sets, which are isomorphic on each special open sets (3.5) and compatible with the natural restrictions. Then considering there sheafifications and using the above isomorphisms, we get isomorphism between the complexes of sheaves, described in (15).
@@
for each $U_{x}$ associated with $x\in X$, where $P_{U_{x}}$ is given by the cannonical projections $pr_{1}:\mathcal{J}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))\cong\widehat{S_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})^{*}}\rightarrow\mathcal{L}(U_{x})^{*}$ and the anti-symmetrization map $Alt_{U_{x}}:\wedge^{\bullet}_{\mathcal{O}_{X}(U_{x})}\mathcal{L}(U_{x})^{*}\rightarrow\mathcal{J}(\mathcal{O}_{X}(U_{x}),\mathcal{L}(U_{x}))^{\otimes^{\bullet}}$ for every special open set $U_{x}$ of $X$ (for smooth manifold we get these quasi-isomorphisms on each open sets). Sheafification of the associated presheaves provides the quasi-isomorphism between the cobar complex $\mathscr{C}ob^{\bullet}(\mathscr{J}(\mathcal{O}_{X},\mathcal{L}))$ and the dual Koszul-Rinehart complex $\bar{\mathscr{K}}^{\bullet}_{\mathcal{O}_{X}}\mathcal{L}^{*}$.
@@
In [5, 6], a dual version of the (twisted) HKR theorem for Lie algebroids [7], utilizing different techniques to handle rich algebraic structures, is considered to study the notion of precalculus up to homotopy in the context of Lie algebroids over ringed sites.
@@
All the previous discussions are applicable to Lie algebroids over Noetherian separated schemes (or schemes of finite types) over a field of characteristic zero (see [18, 3]). Moreover, these results can be extended to the more general notion of Lie algebroids over ringed sites ([7, 6, 5]). It requires more of an algebraic viewpoint rather than a geometric one.